OpenStudy (anonymous):

integral of sec (3x) tan (3x) dx

OpenStudy (rogue):

If the derivative of sec x is sec x tan x, what do you think is the integral of sec (3x) tan (3x)? :P

OpenStudy (anonymous):

Let u = 3x du = 3 dx 1/3 du = dx (1/3)\[\int\limits_{?}^{?}(secu*tanu)du\] (1/3) secu + c (1/3) sec(3x)

OpenStudy (anonymous):

The (1/3) above the integral is supposed to be next to the integral, multiplying it by 1/3.